Hi I am so lost and cant figure this out please help me!!!
Find a rational number, r, that satisfies the following conditions. 3 separate answers.
(a) r^2 < r
(b) r^2 = r
(c) r^2 > r

Answer by mananth(12270)

(a) r^2 < r
Let r = 1/2
(1/2)^2 = 1/4
!/4 < 1/2

(b) r^2 = r
let r=1
(1)^2 = 1
(c) r^2 > r
Let r= 2
2^2=4
4>2